# State-Covering Trajectory Stitching for Diffusion Planners

### v2

@@ 4.4 Offline GCRL with SCoTS-Augmented Data
Although SCoTS is primarily designed for diffusion planners, we additionally evaluate whether trajectories augmented by SCoTS can enhance the performance of existing offline goal-conditioned RL (GCRL) algorithms. Specifically, we retrain widely used offline GCRL algorithms, including GCIQL Kostrikov et al. (2022), CRL Eysenbach et al. (2022), and HIQL Park et al. (2023a), on the SCoTS-augmented dataset. All hyperparameters remain identical to their original implementations. Additionally, we compare our approach with SynthER Lu et al. (2023), which employs an unconditional diffusion model for transition-level data augmentation. Results summarized in Table 2 clearly demonstrate that SCoTS-generated trajectories consistently outperform SynthER and methods trained solely on the original offline datasets, significantly boosting performance across all tested algorithms. This indicates that augmenting data at the trajectory-level with SCoTS, which explicitly considers long-term dynamics and diversity, provides more effective supervision for learning robust trajectory stitching and long-horizon planning capabilities.
@@ 4.5 Ablation Studies
Dynamic $\text{Dynamic MSE}=\|f^{*}(\boldsymbol{s},\boldsymbol{a})-\boldsymbol{s}^{\prime}\|_{2}^{2},$
@@
We employs replanning during a rollout, enabling the agent to recover from failures, such as when the diffusion planner generates unreachable subgoals for the low-level controller. In practice, we set a replanning interval (e.g., every 200 steps); further implementation details are provided in Appendix B. In Table 3, we present an ablation study comparing performance with and without replanning on the PointMaze and AntMaze Stitch datasets from OGBench. SCoTS consistently outperforms Compositional Diffuser (CD) Luo et al. (2025), the best-performing baseline, even without replanning. Additionally, the performance with and without replanning is similar, highlighting the reliability and efficacy of the SCoTS-augmented diffusion planner.
@@ 5 Related Work
Diffusion probabilistic models Sohl-Dickstein et al. (2015); Ho et al. (2020) have emerged as powerful tools for reinforcement learning, especially in offline settings. These models iteratively denoise sampled data from noise, effectively learning gradients of the data distribution Song & Ermon (2019) and demonstrating strong capabilities in modeling complex trajectories. Early work such as Diffuser Janner et al. (2022) employed unconditional diffusion models guided by learned value estimators Dhariwal & Nichol (2021). Subsequent methods like Decision Diffuser Ajay et al. (2023) and AdaptDiffuser Liang et al. (2023) introduced classifier-free guidance and progressive fine-tuning. Recent advancements further leveraged hierarchical structures Chen et al. (2024c); Li et al. (2023), multi-task conditioning Ni et al. (2023); He et al. (2023); Dong et al. (2024), and multi-agent setups Zhu et al. (2023). Additionally, diffusion planners have explored integration with tree search methods Yoon et al. (2025), refined trajectory sampling techniques Lee et al. (2023); Feng et al. (2024); Lee & Choi (2025), and investigated critical design choices to improve robustness Lu et al. (2025). Despite these advances, diffusion planners still fundamentally depend on the quality and diversity of the offline training datasets, limiting their ability to generate coherent and feasible long-horizon plans beyond their training distribution. Recent approaches such as Generative Skill Chaining (GSC) Mishra et al. (2023) and Compositional Diffuser Luo et al. (2025) address this by composing short segments at test time into long-horizon trajectories. Our work presents an orthogonal solution by directly augmenting the offline dataset itself, significantly enhancing the capability of diffusion planners to generalize to diverse and substantially longer trajectories.
@@
Temporal distance has been widely adopted as a structural inductive bias in various reinforcement learning (RL) paradigms, including imitation learning (Sermanet et al., 2018), unsupervised skill discovery (Hartikainen et al., 2019; Park et al., 2023b, 2024b), goal-conditioned RL (Durugkar et al., 2021; Eysenbach et al., 2022; Wang et al., 2023; Bae et al., 2024), and curriculum learning (Zhang et al., 2020; Kim et al., 2023). Recent methods such as METRA Park et al. (2023b), QRL Wang et al. (2023), HILP Park et al. (2024b), and TLDR Bae et al. (2024) particularly focus on learning temporal distance-preserving representations to facilitate diverse skill discovery or efficient goal-reaching behaviors. Distinct from prior methods, our SCoTS framework explicitly leverages temporal distance-preserving representations to identify temporally viable trajectory segments for stitching. This allows systematic synthesis of extended, diverse, and dynamically consistent trajectories, significantly augmenting offline datasets and improving long-horizon generalization for diffusion-based planners.
@@ Appendix A Details of Datasets
We evaluate our method on the OGBench benchmark Park et al. (2024a)https://github.com/seohongpark/ogbench. Since our primary goal is to assess trajectory stitching capability and long-horizon reasoning, we specifically utilize the Stitch and Explore datasets. As shown in Figure 7, the Stitch dataset is explicitly designed to challenge trajectory stitching ability, comprising short, goal-reaching trajectories limited to a maximum length of four cell units. Consequently, agents must effectively stitch together multiple short segments (up to eight) to successfully complete long-horizon tasks. In contrast, the Explore dataset is designed to test navigation skills learned from extensive yet low-quality trajectories. These trajectories are generated by commanding a low-level policy with random movement directions re-sampled every ten steps, along with significant action noise. Each demonstration trajectory typically spans only two to three blocks, resulting in noisy and clustered paths that pose additional challenges for evaluating the ability to learn effective policies from highly suboptimal data.
@@ Appendix B Implementation Details
Details of the low-level controller.
A key challenge in diffusion-based planning is balancing global trajectory coherence with effective low-level control in high-dimensional state-action spaces Chen et al. (2024a, b); Yoon et al. (2025). Previous approaches, such as PlanDQ Chen et al. (2024b) and MCTD Yoon et al. (2025), address this issue by integrating high-level diffusion planners with separately trained low-level controllers. Similarly, we adopt a hierarchical strategy, where the diffusion planner generates plans based primarily on compact, lower-dimensional state representations (e.g., positions of the agent itself), delegating the fine-grained, low-level action execution to a dedicated low-level controller. In our experiments, we specifically employ GCIQL Kostrikov et al. (2022) as the learned low-level policy in the PointMaze environments and CRL Eysenbach et al. (2022) in the AntMaze environments. A detailed visualization of generated subgoals and their corresponding execution rollouts can be seen in Figure 9. Furthermore, an ablation study examining the impact of the horizon length of the low-level controller is presented in Figure 6.
Implementation details for SCoTS and diffusion planning.
Practical implementation of temporal distance-preserving search.
Our SCoTS framework relies on a learned latent space $\mathcal{Z}$ where the $L_{2}$ distance, $\|\phi(\boldsymbol{s})-\phi(\boldsymbol{g})\|_{2}$, approximates the optimal temporal distance $d^{*}(\boldsymbol{s},\boldsymbol{g})$ between states (as detailed in Section 3.1). A critical step in SCoTS is the efficient identification of suitable candidate trajectory segments from a large offline dataset $\mathcal{D}$. This requires a fast nearest neighbor search mechanism within the learned latent space $\mathcal{Z}$. To achieve this, we employ an Inverted File (IVF) index from the Faiss library Douze et al. (2024), which is specifically designed for large-scale similarity searches.
@@
Learning Rate $3\times 10^{-4}$
@@
Learning Rate $2\times 10^{-4}$
Weight Decay $1\times 10^{-5}$
@@
Learning Rate $2\times 10^{-4}$
Weight Decay $1\times 10^{-5}$
@@ Appendix C Additional Results
We train temporal distance-preserving latent representations with dimension 32 across all environments. To visualize these learned representations, we apply a $t$-distributed stochastic neighbor embedding (t-SNE) to project the 32-dimensional latent vectors onto a 2-dimensional plane, as shown in Figure 8. Recall from Equation 6 that we parameterize a goal-conditioned value function $V(\boldsymbol{s},\boldsymbol{g})$ following Park et al. (2024b):
$V(\boldsymbol{s},\boldsymbol{g})\coloneqq-||\phi(\boldsymbol{s})-\phi(\boldsymbol{g})||_{2},$ (13)
which approximates the optimal goal-conditioned value function, defined as the maximum possible return (cumulative sum of rewards) for sparse-reward settings. Specifically, an agent receives a reward of 00 if the $l_{2}$ distance between states $\boldsymbol{s}$ and $\boldsymbol{g}$ is within a small threshold $\delta_{g}$, and $-1$ otherwise. The embedding function $\phi$ is trained using a temporal-difference objective inspired by implicit Q-learning Kostrikov et al. (2022) on the offline dataset $\mathcal{D}$. As illustrated in Figure 8, the learned representations effectively capture the temporal proximity between states, resulting in latent spaces where states that are temporally close in the environment are also clustered closely in the embedding space.

### v3

@@ 4.4 Offline GCRL with SCoTS-Augmented Data
Although SCoTS is primarily designed for diffusion planners, we additionally evaluate whether trajectories augmented by SCoTS can enhance the performance of existing offline goal-conditioned RL (GCRL) algorithms. Specifically, we retrain widely used offline GCRL algorithms, including GCIQL Kostrikov et al. (2022), CRL Eysenbach et al. (2022), and HIQL Park et al. (2023), on the SCoTS-augmented dataset. All hyperparameters remain identical to their original implementations. Additionally, we compare our approach with SynthER Lu et al. (2023), which employs an unconditional diffusion model for transition-level data augmentation. Results summarized in Table 2 clearly demonstrate that SCoTS-generated trajectories consistently outperform SynthER and methods trained solely on the original offline datasets, significantly boosting performance across all tested algorithms. This indicates that augmenting data at the trajectory-level with SCoTS, which explicitly considers long-term dynamics and diversity, provides more effective supervision for learning robust trajectory stitching and long-horizon planning capabilities.
@@ 4.5 Ablation Studies
$\text{Dynamic MSE}=\|f^{*}(\boldsymbol{s},\boldsymbol{a})-\boldsymbol{s}^{\prime}\|_{2}^{2},$
@@
We employ replanning during a rollout, enabling the agent to recover from failures, such as when the diffusion planner generates unreachable subgoals for the low-level controller. In practice, we set a replanning interval (e.g., every 200 steps); further implementation details are provided in Appendix B. In Table 3, we present an ablation study comparing performance with and without replanning on the PointMaze and AntMaze Stitch datasets from OGBench. SCoTS consistently outperforms Compositional Diffuser (CD) Luo et al. (2025), the best-performing baseline, even without replanning. Additionally, the performance with and without replanning is similar, highlighting the reliability and efficacy of the SCoTS-augmented diffusion planner.
@@ 5 Related Work
Diffusion probabilistic models Sohl-Dickstein et al. (2015); Ho et al. (2020) have emerged as powerful tools for reinforcement learning, especially in offline settings. These models iteratively denoise sampled data from noise, effectively learning gradients of the data distribution Song & Ermon (2019) and demonstrating strong capabilities in modeling complex trajectories. Early work such as Diffuser Janner et al. (2022) employed unconditional diffusion models guided by learned value estimators Dhariwal & Nichol (2021). Subsequent methods like Decision Diffuser Ajay et al. (2023) and AdaptDiffuser Liang et al. (2023) introduced classifier-free guidance and progressive fine-tuning. Recent advancements further leveraged hierarchical structures Chen et al. (2024c); Li et al. (2023), multi-task conditioning Ni et al. (2023); He et al. (2023); Dong et al. (2024), and multi-agent setups Zhu et al. (2023). Additionally, diffusion planners have explored integration with tree search methods Yoon et al. (2025a, b), refined trajectory sampling techniques Lee et al. (2023b); Feng et al. (2024); Lee & Choi (2025), and investigated critical design choices to improve robustness Lu et al. (2025). Despite these advances, diffusion planners still fundamentally depend on the quality and diversity of the offline training datasets, limiting their ability to generate coherent and feasible long-horizon plans beyond their training distribution. Recent approaches such as Generative Skill Chaining (GSC) Mishra et al. (2023) and Compositional Diffuser Luo et al. (2025) address this by composing short segments at test time into long-horizon trajectories. Our work presents an orthogonal solution by directly augmenting the offline dataset itself, significantly enhancing the capability of diffusion planners to generalize to diverse and substantially longer trajectories.
@@
Temporal distance has been widely adopted as a structural inductive bias in various reinforcement learning (RL) paradigms, including imitation learning (Sermanet et al., 2018), unsupervised skill discovery (Hartikainen et al., 2019; Park et al., 2024b, a), goal-conditioned RL (Durugkar et al., 2021; Eysenbach et al., 2022; Wang et al., 2023; Bae et al., 2024), and curriculum learning (Zhang et al., 2020; Kim et al., 2023). Recent methods such as METRA Park et al. (2024b), QRL Wang et al. (2023), HILP Park et al. (2024a), and TLDR Bae et al. (2024) particularly focus on learning temporal distance-preserving representations to facilitate diverse skill discovery or efficient goal-reaching behaviors. Distinct from prior methods, our SCoTS framework explicitly leverages temporal distance-preserving representations to identify temporally viable trajectory segments for stitching. This allows systematic synthesis of extended, diverse, and dynamically consistent trajectories, significantly augmenting offline datasets and improving long-horizon generalization for diffusion-based planners.
@@ 6 Conclusion
Impact Statement.
This paper advances the field of diffusion-based planning by introducing a novel trajectory augmentation method, enhancing long-horizon reasoning capabilities in reinforcement learning. While we do not identify direct negative societal impacts stemming from this research, practitioners are encouraged to carefully assess real-world implications, particularly regarding safety and reliability, prior to deploying this method in practical scenarios.
Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.2019-0-00075, Artificial Intelligence Graduate School Program (KAIST); No. 2022-0-00984, Development of Artificial Intelligence Technology for Personalized Plug-and-Play Explanation and Verification of Explanation; No. RS-2024-00457882, AI Research Hub Project; No.RS-2022-II220184, Development and Study of AI Technologies to Inexpensively Conform to Evolving Policy on Ethics), and by the InnoCORE program of the Ministry of Science and ICT (N10250156).
@@ Appendix A Details of Datasets
We evaluate our method on the OGBench benchmark Park et al. (2025) [^1] https://github.com/seohongpark/ogbench. Since our primary goal is to assess trajectory stitching capability and long-horizon reasoning, we specifically utilize the Stitch and Explore datasets. As shown in Figure 7, the Stitch dataset is explicitly designed to challenge trajectory stitching ability, comprising short, goal-reaching trajectories limited to a maximum length of four cell units. Consequently, agents must effectively stitch together multiple short segments (up to eight) to successfully complete long-horizon tasks. In contrast, the Explore dataset is designed to test navigation skills learned from extensive yet low-quality trajectories. These trajectories are generated by commanding a low-level policy with random movement directions re-sampled every ten steps, along with significant action noise. Each demonstration trajectory typically spans only two to three blocks, resulting in noisy and clustered paths that pose additional challenges for evaluating the ability to learn effective policies from highly suboptimal data.
@@ Appendix B Implementation Details
Learning Rate $3\times 10^{-4}$ -
@@
Learning Rate $2\times 10^{-4}$ -
Weight Decay $1\times 10^{-5}$ -
@@
Learning Rate $2\times 10^{-4}$ -
Weight Decay $1\times 10^{-5}$ -
@@
Details of the low-level controller.
A key challenge in diffusion-based planning is balancing global trajectory coherence with effective low-level control in high-dimensional state-action spaces Chen et al. (2024a, b); Yoon et al. (2025a). Previous approaches, such as PlanDQ Chen et al. (2024b) and MCTD Yoon et al. (2025a), address this issue by integrating high-level diffusion planners with separately trained low-level controllers. Similarly, we adopt a hierarchical strategy, where the diffusion planner generates plans based primarily on compact, lower-dimensional state representations (e.g., positions of the agent itself), delegating the fine-grained, low-level action execution to a dedicated low-level controller. In our experiments, we specifically employ GCIQL Kostrikov et al. (2022) as the learned low-level policy in the PointMaze environments and CRL Eysenbach et al. (2022) in the AntMaze environments. A detailed visualization of generated subgoals and their corresponding execution rollouts can be seen in Figure 9. Furthermore, an ablation study examining the impact of the horizon length of the low-level controller is presented in Figure 6.
Implementation details for SCoTS and diffusion planning.
Practical implementation of temporal distance-preserving search.
Our SCoTS framework relies on a learned latent space $\mathcal{Z}$ where the $L_{2}$ distance, $\|\phi(\boldsymbol{s})-\phi(\boldsymbol{g})\|_{2}$, approximates the optimal temporal distance $d^{*}(\boldsymbol{s},\boldsymbol{g})$ between states (as detailed in Section˜3.1). A critical step in SCoTS is the efficient identification of suitable candidate trajectory segments from a large offline dataset $\mathcal{D}$. This requires a fast nearest neighbor search mechanism within the learned latent space $\mathcal{Z}$. To achieve this, we employ an Inverted File (IVF) index from the Faiss library Douze et al. (2024), which is specifically designed for large-scale similarity searches.
@@ Appendix C Additional Results
Ablation study on temporal distance-preserving embedding.
Table 6: Ablation on the temporal distance-preserving embedding. Success rates on OGBench PointMaze and AntMaze Stitch datasets. ✗ indicates stitching guided by raw state-space distances (w/o temporal embedding), while ✓ indicates guidance from learned temporal embedding.
Env Size SCoTS
[…]
Average 49.3 95.0
To isolate the contribution of our learned latent space, we compare the performance of the HD Chen et al. (2024c) planner trained on SCoTS-augmented data where the stitching process was guided by either our temporal distance-preserving embedding or raw state space. As shown in Table 6, temporal distance-preserving representation is crucial for effective stitching. This is especially true in high-dimensional environments like AntMaze, where a small Euclidean distance in the raw state space (e.g., between two similar joint configurations) does not guarantee reachability. Even in Pointmaze, where state-space distance is more intuitive, the learned compact latent space provides a better-structured representation for stitching temporally extended trajectories.
Ablation study on diffusion-based stitching refinement.
Table 7: Ablation on diffusion-based stitching refinement. Success rates on AntMaze Stitch datasets for HD Chen et al. (2024c) and HIQL Park et al. (2023). ✗ indicates training on SCoTS data generated without refinement, while ✓ indicates training on data generated with refinement.
Env Size HD HIQL
[…]
Average 69.0 90.0 31.5 73.0
To demonstrate the importance of the refinement step, we compare the performance of both a diffusion planner (HD) and a GCRL algorithm (HIQL) trained on SCoTS data generated with and without the diffusion-based refinement. The results, summarized in Table 7, reveal the critical importance of this component. Without refinement, the connection points between stitched segments can suffer from large dynamic inconsistency errors, as illustrated in Figure 6. This performance degradation is particularly significant when training GCRL algorithms like HIQL, which are highly sensitive to the dynamic validity of the training transitions.
Sensitivity analysis on the novelty weight $\beta$.
As illustrated in Figure 3, using only the progress score ($\beta=0$), which quantifies the alignment with a latent exploration direction, is effective at generating long, temporally-extended trajectories with clear directional distinctions. However, relying solely on this directional guidance can be suboptimal. The learned temporal distance-preserving embedding $\phi$ is a powerful but imperfect approximation of the true environment topology. Due to embedding errors, the geometry of the latent space can be distorted. Consequently, a straight line between two latent states, $\phi(s)$ and $\phi(s^{\prime})$, may not correspond to a feasible path in the underlying MDP. Therefore, exploratory detours are often necessary to find a valid temporally extended path. This is precisely where the novelty score becomes critical. It encourages these exploratory detours by rewarding the selection of segments that lead to less-visited states. This allows the agent to navigate around the imperfections and distortions in the learned latent space, discovering feasible and often more effective paths. To demonstrate this complementary effect quantitatively, we conducted an additional ablation study on 2 OGBench tasks. The table below compares the performance of the HD planner trained with SCoTS data generated using different novelty weights $\beta$ (5 seeds). The results, summarized in Table 8, suggest that while directional stitching alone ($\beta=0$) provides a notable performance improvement, introducing and balancing it with the novelty score ($\beta>0$) yields substantial further gains. This effect is especially pronounced on the more complex Giant task.
Table 8: Sensitivity to novelty weight $\beta$.
Env Type Size HD $\beta$=0 $\beta$=2 $\beta$=4 $\beta$=8
[…]
Sensitivity analysis on sub-trajectory length $H$.
As shown in Table 9, performance on the Stitch dataset is relatively robust to the choice of $H$. However, on the Explore dataset, shorter segments perform best. We hypothesize this is because the Explore dataset consists of low-quality, noisy trajectories. Using shorter segments allows SCoTS to be more selective, finding and connecting the temporally-extended parts of trajectories.
Table 9: Sensitivity to sub-trajectory length $H$.
Env Type Size HH=26 HH=52 HH=104
[…]
Sensitivity analysis on the number of retrieved segments $K$.
Table 10 shows that performance is robust as long as $K$ is not too small. A very small $K$ limits the diversity of candidate segments, hindering the effectiveness of our progress and novelty-based selection. While performance is stable for larger $K$, we expect an excessively large $K$ could eventually degrade performance by increasing the chance of retrieving dynamically inconsistent segments.
Table 10: Sensitivity to number of retrieved segments $K$.
Env Type Size KK=3 KK=10 KK=20
[…]
Evaluation on dense-reward offline RL benchmarks.
While our paper focuses on tasks where extrinsic rewards are absent, the SCoTS framework is not limited to goal-conditioned tasks. To demonstrate its applicability to general, dense-reward offline RL, we conducted additional experiments on four standard MuJoCo locomotion benchmarks from D4RL Kumar et al. (2020), utilizing suboptimal medium and medium-replay datasets. It is important to clarify that our use of reward-free refers to the stitching mechanism itself, which is guided by the intrinsic temporal structure of the data rather than extrinsic task rewards. This allows the core data augmentation to be task-agnostic.
For a fair comparison, we used the same Diffuser architecture and planning horizon (32) as described in the original paper Janner et al. (2022). We upsample the original data to 2M samples using the same SCoTS procedure, with sub-trajectory lengths set to 4. Additionally, to label these new trajectories with rewards, we additionally trained a reward model, $r_{t}=f_{\omega}(s_{t},a_{t})$, on the original offline dataset. As shown in Table 11, training on SCoTS-augmented data consistently improves the performance of the original Diffuser. Even without architectural or sampling strategy changes, by stitching sub-trajectories in a way that is temporally extended and exploratory, SCoTS generates novel behaviors (e.g., trajectories that travel further) that are not present in the original suboptimal dataset.
Table 11: Performance on D4RL benchmarks. We report normalized average returns of the Diffuser and Diffuser trained on SCoTS-augmented data. Results are averaged over 50 planning seeds.
Dataset Diffuser Diffuser + SCoTS (ours)
[…]
Average 79.5 87.2
Evaluation on noisy and partially observable navigation.
To evaluate the robustness of SCoTS on noisy and low-quality data beyond the controlled OGBench settings, we conducted additional experiments using a mobile robot navigation task in realistic, large-scale environments (a 35m x 39m corridor and a 50m x 110m building), following the setup from HRL Lee et al. (2023a). We simulated low-quality data by generating short, disconnected segments. Specifically, trajectories were collected by randomly selecting start and goal locations within short ranges (5m) using HRL Lee et al. (2023a). To handle the noisy raw 2D LiDAR measurements (512 rays over 360-degree), we converted these scans into 64x64 occupancy grid images. These images were then encoded using a pre-trained $\beta$-VAE Higgins et al. (2017) into a compact 8-dimensional latent representation, providing structured, low-dimensional state inputs for the agent. Subsequently, the SCoTS stitching method was applied to augment this fragmented dataset. The results, shown in Table 12, clearly indicate that SCoTS effectively enhances performance, even with highly fragmented and noisy datasets, demonstrating its robustness and practical applicability to robot navigation tasks.
Table 12: Performance on a mobile robot navigation task. Goal-reaching success rates are reported over 100 episodes for the HIQL Park et al. (2023) trained with and without SCoTS augmentation.
Environment HIQL HIQL + SCoTS (ours)
[…]
@@
We train temporal distance-preserving latent representations with dimension 32 across all environments. To visualize these learned representations, we apply a $t$-distributed stochastic neighbor embedding (t-SNE) to project the 32-dimensional latent vectors onto a 2-dimensional plane, as shown in Figure 8. Recall from Equation 6 that we parameterize a goal-conditioned value function $V(\boldsymbol{s},\boldsymbol{g})$ following Park et al. (2024a):
[…]
which approximates the optimal goal-conditioned value function, defined as the maximum possible return (cumulative sum of rewards) for sparse-reward settings. Specifically, an agent receives a reward of 0 if the $l_{2}$ distance between states $\boldsymbol{s}$ and $\boldsymbol{g}$ is within a small threshold $\delta_{g}$, and 1-1 otherwise. The embedding function $\phi$ is trained using a temporal-difference objective inspired by implicit Q-learning Kostrikov et al. (2022) on the offline dataset $\mathcal{D}$. As illustrated in Figure 8, the learned representations effectively capture the temporal proximity between states, resulting in latent spaces where states that are temporally close in the environment are also clustered closely in the embedding space.